Using the values for the heat of fusion, specific heat of water, and/or heat of vaporization, calculate the amount of heat energy in each of the following.

calories need to melt 65.0 g of ice at 0 ∘C and to warm the liquid to 65 ∘C
1
Expert's answer
2020-09-15T13:51:55-0400

Amount of heat energy = mL +mc"\Delta"T ="65\times80+65\times1\times65=5200+4225=9425 \ cal"